### v2

# Echo: Decoupling Inference and Training for Large-Scale RL Alignment on Heterogeneous Swarms

Abstract
Modern RL-based post-training for large language models (LLMs) co-locate trajectory sampling and policy optimisation on the same GPU cluster, forcing the system to switch between inference and training workloads. This serial context switching violates the single-program–multiple-data (SPMD) assumption underlying today’s distributed training systems. We present Echo, the RL system that cleanly decouples these two phases across heterogeneous “inference” and “training” swarms while preserving statistical efficiency. Echo introduces two lightweight synchronization protocols: a sequential pull mode that refreshes policy weights according to API call for minimal bias, and an asynchronous push–pull mode that streams version-tagged rollouts through a replay buffer to maximise hardware utilisation. Training three representative RL workloads with Qwen3-4B, Qwen2.5-7B, and Qwen3-32B on a geographically distributed cluster, Echo matches a fully co-located Verl baseline in convergence speed and final reward while off-loading trajectory generation to commodity edge hardware. These promising results demonstrate that large-scale RL for LLMs could achieve datacentre-grade performance using decentralised, heterogeneous resources.
[…]
1 Introduction
Reinforcement learning (RL) has become the de facto recipe for aligning billion-parameter language models with human preferences, such as RLHFouyang2022training , RLAIFlee2023rlaif , DPOrafailov2023direct and their variants. All of these algorithms share a multi-program, multi-data (MPMD) execution pattern: (i) an inference phase that generates trajectories under the current policy, followed by (ii) a training phase that converts those trajectories into gradient updates. State-of-the-art systems, e.g., DeepSpeed RLHFyao2023deepspeed and Verlsheng2024hybridflow , run the two phases on the same high-bandwidth cluster, implicitly synchronising through global barriers and expensive NVLink / InfiniBand fabrics. Since trajectory sampling and policy optimization have very different hardware footprints, running both phases on a single hardware pool forces the cluster to serially contextual-switch between two different workloads. This back-and-forth violates the single-program–multiple-data (SPMD) assumption that underpins most large-scale training systems, hindering utilisation and scalability.
The natural remedy is to split trajectory generation and policy optimization onto distinct, heterogeneous swarms. However, naive decoupling may break the implicit guarantees provided by co-location: rollouts may be generated by stale policies; learners may optimise with out-of-date data; and, as a result, the entire training can diverge.
In this paper, we introduce Echo, the RL system that solves the above problem through a pair of synchronization mechanisms, as shown in Figure 1. The sequential mode has trainers pull rollouts via an API, refreshing policy weights beforehand to mirror classic PPO and minimise bias. The asynchronous mode lets samplers push tagged rollouts into a replay buffer while a coordinator keeps tracking version drift, maximising device utilisation.
On the inference side, we build upon Parallax, a fully decentralised pipeline-parallel engine developped by Gradient Network that converts geographically dispersed consumer GPUs and Apple-Silicon devices into a single, high-throughput sampler via dynamic KV-cache management and continuous batching. On the training side we adopt the community-standard Verl stack, adding robust support for LoRA-style parameter-efficient finetuning.
We train three typical RL workloads from scratch on Qwen3-4B, Qwen2.5-7B, and Qwen3-32B—using both Echo and a fully co-located Verl baseline. In every case, Echo matches the baseline’s convergence speed and final reward while off-loading trajectory generation to a pool of heterogeneous edge devices. These results demonstrate that large-scale RL can exploit decentralized, commodity hardware without sacrificing training efficiency.
Contributions.
  1. 1.
    We formulate the information-synchronisation challenge arising from decoupling RL trajectory generation and policy optimisation.
  2. 2.
    We design two complementary protocols, i.e., Sequential and Asynchronous, that bound policy lag without datacentre-class interconnects.
  3. 3.
    We implement Echo, featuring the Parallax inference engine and an enhanced Verl trainer with LoRA support.
  4. 4.
    Our extensive experiments show that a heterogeneous, geographically distributed cluster can train LLMs via RL at parity with high-end datacenter baselines.
2 System Design
2.1 System Overview
Conventional RL stacks co-locate trajectory generation and policy updates on the same high-bandwidth cluster, implicitly relying on barrier-style synchronisation to guarantee that every gradient step consumes up-to-date data. Echo deliberately splits these two phases onto distinct, heterogeneous swarms—an inference swarm that samples trajectories and a training swarm that optimises the policy—so that each can run on the hardware best suited to its workload. The price of this flexibility is a non-trivial information synchronisation problem: trajectory freshness and weight staleness must be controlled explicitly; otherwise, the decoupled system may diverge or converge more slowly than monolithic baselines. Tackling this challenge requires a careful analysis of the data-dependency chain between policy parameters, sampled trajectories, and gradient updates.
Figure 1: Echo architecture and the two synchronisation mechanisms between the training and inference swarms.
As shown in Figure 1, the Echo system is composed of two complementary protocols focusing on different optimisation goals:
  • Sequential mechanism (accuracy-centric). The training swarm pulls trajectories on demand via an API, and the inference peer first checks the version of its local weights before producing the rollouts. Therefore, the version of the policy weights used to generate rollouts can be precisely controlled. This path delivers the lowest bias and therefore the best statistical accuracy.
  • Asynchronous mechanism (efficiency-centric). The inference swarm streams rollouts continuously to a shared rollout buffer, tagging each sample with the parameter version. The training swarm consumes these version-stamped batches at its own pace and requests a weight refresh only when the observed version drift exceeds a threshold. This path maximises hardware utilisation and is preferred when abundant, under-used inference capacity is available.
Both protocols adopt model-snapshot buffer that stores the model checkpoints to be synchronized. Together Echo let users trade a small amount of statistical bias for a large gain in throughput, or vice versa, without altering the rest of the system.
Beyond cross-swarm synchronisation, Echo embeds state-of-the-art optimisations inside each swarm. In the Inference side, We build upon Parallax, a comprehensive distributed inference framework that runs LLMs seamlessly across heterogeneous devices—from Nvidia GPUs to Apple-Silicon laptops. The training side extends the community-standard Verl RL stack, preserving broad algorithmic compatibility (PPO, GRPO, DPO, …) while adding first-class support for parameter-efficient adapters such as LoRA.
The details of each component of Echo are described in the following sections.
2.2 Synchronization Mechanism
Decoupling rollout generation from gradient calculation transforms the classical, single-loop RL pipeline into two semi-independent yet mutually-constrained swarms. In this section, we describe how (i) trajectories flow upward from the inference swarm to the training swarm and (ii) policy snapshots flow downward in the opposite direction, while keeping the correctness of the end-to-end training computation.
2.2.1 Rollouts
Our framework exposes two complementary data paths from the inference swarm back to the trainers.
1) Inference-as-an-API (trainer-driven, pull-based).
In this mode the inference swarm surfaces a trajectory-generation API. The training swarm invokes this endpoint whenever it needs fresh data, thereby explicitly deciding what to sample, when to sample, and whether the sampler should refresh its local model weights before serving the request.
  • API contract. The endpoint follows an OpenAI-compatible JSON schema so that existing tooling (e.g. OpenAI Python SDK) can be reused:
    POST /v1/rl/trajectories
      "model": "ECHO-Infer",
      "param_version": "theta_e",   // trainer’s current weights
      "prompts": [p_1, , p_k],
      "algo": "PPO"                 // or PPOvariant identifier
      "param_version": "theta_e",   // echoed back for traceability
      "trajectories": [
        { "tokens": [...],
          "actions": [...],
          "logprobs": [...],
          "values": [...],
          "rewards": [...],
          "terminal": true },
    Every trajectory yields exactly the $(s,a,\log\pi_{\theta}(a|s),v_{\theta}(s),r)$ fields required by PPO and its popular variants.
  • Call logic. From the trainer’s perspective the call is blocking: optimisation does not proceed until the requested batch returns. The simplest policy is to invoke the API once per optimisation step, but more sophisticated schedulers can batch multiple steps or interleave asynchronous prefetchers. On the inference side the request handler (i) compares its local weights against the caller’s param_version; if the version gap exceeds a pre-configured threshold it reloads the newer weights, otherwise it continues with the current checkpoint; (ii) executes the rollout script to produce $k$ trajectories; (iii) streams the batch back to the caller.
  • Applicable scenarios. This pull-based path is ideal when the algorithm demands fine-grained, runtime control over both the data distribution and the policy version—e.g. active-learning curricula, on-policy off-diagonal exploration, or methods with low tolerance to policy staleness. Because the semantics match those of a single-process PPO loop, researchers can port centrally-designed RL code without rewriting dataflow logic. The blocking nature of the call makes it best suited to short rollouts or compute-rich inference swarms where latency is dominated by a single forward pass rather than network transfers.
2) Replay-Buffer Streaming (asynchronous, push–pull).
In this regime the inference and training swarms progress concurrently rather than in a strict master-worker loop. The inference swarm continuously produces trajectories with its current policy and pushes them into a distributed replay buffer, while the training swarm pulls mini-batches from that buffer at its own pace. No direct dependency exists between the two loops except for co-ordinated weight synchronisations.
  • Overlapping the rollout generation and the training. Once optimisation starts, every inference node autonomously samples the environment with its resident policy, annotates each rollout with a param_version tag, and appends the resulting batch to the replay buffer. Meanwhile the training swarm streams mini-batches from the buffer and updates the policy in parallel, so rollout generation and gradient computation overlap in time, amortising end-to-end latency.
  • Coordinator. Because the strong synchronization is intentionally broken, a lightweight coordinator orchestrates two critical invariants. First, it balances throughput by aligning the inference batch size to an integer multiple of the trainer’s mini-batch size, mitigating “batch fragmentation” in the replay buffer. Second, it monitors the version skew between the policy used to produce a batch and the one used to consume it; when the skew exceeds a configurable threshold, it broadcasts a sync_weight to both sides.
  • Applicable scenarios. This path suits algorithms that tolerate limited policy staleness—e.g. experience replay variants of PPO or off-policy critics—because generated trajectories may lag one or more updates behind the current actor. It is particularly advantageous when individual trajectories involve long-horizon or multi-turn inference, where overlapping compute can mask inference latency and yield a substantial utilisation improvement.
2.2.2 Model Weights
In a centralised PPO loop, the same policy is used for rollout generation and gradient updates, so actor–learner staleness is naturally bounded. After decoupling the two swarms, however, the convergence efficiency highly depends on keeping their weight versions sufficiently aligned. Whenever synchronisation is required, the trainer serialises its current checkpoint, uploads it to model-snapshot buffer, and the inference swarm fetches and activates the new policy.
  • API-triggered synchronisation. In the Inference-as-an-API path the trainer embeds its latest param_version in every request. Upon receipt the inference node compares this tag with the version of its resident weights; if the numeric or hash distance exceeds a pre-configured threshold, it starts to download the new checkpoint and reloads the latest policy, and only then serves the rollout request.
  • Coordinator-triggered synchronization (replay-buffer mode). For the asynchronous replay-buffer path, the coordinator continually tracks two counters: the learner’s global update step $t_{\text{train}}$ and the current update step used by the inference swarm $t_{\text{infer}}$. When $t_{\text{train}}-t_{\text{infer}}>\Delta_{\max}$, the coordinator broadcasts a sync_weight command. This scheme amortises transfer cost over many trajectories while bounding the maximum policy lag to $\Delta_{\max}$. Note that pre-defining a max staleness degree is just the simplest strategy. By leveraging the Coordinator, more complicated synchronization algorithms can be easily integrated into Echo, such as comparing the distribution distance of the model output, or the statistic features of two models, and so on. We leave this in the future work.
2.3 Fully Decentralized Inference Engine - Parallax
Parallax links a pool of geographically dispersed, consumer-grade accelerators, like 40/50-series GPUs, Apple M-series, and so on, into a single pipeline-parallel sampler. Unlike datacenter engines that rely on RDMA fabrics and homogeneous hardware, Parallax assumes nothing beyond commodity Ethernet and heterogeneous devices.
Layer Allocation.
Here we take a simple yet effective algorithm to allocate the model layers among devices. During initialisation the inference engine probes every node for its peak FLOPS $f_{\text{peak}}$} and free memory $m_{\text{free}}$}. It then partitions the $L$ transformer layers into $S$ static stages $\{\mathcal{L}_{1},\dots,\mathcal{L}_{S}\}$ that satisfy
$\forall s,\quad T_{\mathrm{comp}}(\mathcal{L}_{s})\;\approx\;\frac{1}{S}\sum_{j=1}^{L}T_{\mathrm{comp}}(\ell_{j}),$
so each stage is expected to finish a forward step in roughly the same time. Because this balancing happens once before roll-outs start, the runtime needs no layer migration, credit rebalancing, or token stealing—keeping the wire protocol thin enough for ordinary 10–25 GbE links.
2.4 Training with Algorithm Diversity and Hardware Efficiency
The Echo training swarm consumes trajectory batches, applies a chosen RL algorithm, and performs gradient updates on the policy parameters. Large-language models impose substantial compute and memory pressure, so the training cluster is provisioned with high-bandwidth GPU nodes (e.g. NVLink-connected A100/H100 pods) and exploits a hybrid of data parallelism, pipeline parallelism, and tensor parallelism to scale both model and batch size efficiently. By decoupling training from rollout generation we can separately tune parallel strategy, memory budget, and GPU topology on each side—freeing the training swarm to pursue maximal throughput without constraining the latency-oriented inference swarm.
1) RL Algorithm Diversity.
LLM-centric reinforcement learning is a rapidly evolving research area, with continual proposals for new credit-assignment schemes, preference optimisers, and reward models. Compatibility with the SOTA algorithms is therefore an important design principle. Our framework inherits the modular trainer stack of Verl, one of the community’s most widely adopted RL libraries for language models. Because Verl factorises its optimiser, advantage estimator, and reward inference into interchangeable components, any algorithm implemented in Verl, such as vanilla PPO, KL-constrained PPO, GRPO, DPO, or emerging variants, can be integrated into Echo with negligible glue code; the only requirement is that the optimiser exposes a step() interface consuming mini-batches drawn from the shared buffer.
2) Parameter-efficient Training.
Low-rank adaptation (LoRA) has become an indispensable technique for LLM post-training, reducing effective parameter count by two orders of magnitude and slashing wall-clock time as well as energy cost. In a decentralised setting these savings translate directly into lower checkpoint sizes and thus smaller over-the-wire synchronisation cost. While the upstream Verl repository offers only experimental LoRA support, we contribute a production-ready extension that transparently wraps FSDP and Megatron back-ends.
3 Evaluation
(a) Sokoban w/ Qwen3-4B
(b) Math w/ Qwen2.5-7B
(c) K&K w/ Qwen3-32B
Figure 2: Training curves on the three tasks. Echo always matches the centralized baseline in convergence speed and final return.
The central question we seek to answer by experiments is simple yet crucial: Can a decoupled architecture—where trajectory generation and policy optimisation run on separate, heterogeneous swarms—match the training speed and final performance of a conventional, fully co-located RL stack?
To this end, we reproduce a suite of canonical RL workloads that have well-established learning curves under a datacentre setup and re-train them under Echo.
3.1 Experiment Setup
We conduct experiments on three representative tasks using three model scales from the Qwen series. The primary objective is to compare the performance of the Verl and Echo frameworks across different models and tasks.
Models.
We utilize the open-source Qwen series models as base models for training: Qwen3-4B, Qwen2.5-7B and Qwen3-32B.
Tasks and Datasets.
Sokoban junghanns2001sokoban : We utilize the Sokoban puzzle (Figure 3) to study multi-turn agent interaction in environments requiring irreversible long-term planning. In Sokoban, the agent must push boxes to designated goals within a grid world under constrained steps. Crucially, boxes can only be pushed, not pulled back, demanding significant foresight to avoid irreversible dead ends. We implement this using a Gymbrockman2016openaigym environment configured with a 6 × 6 grid and 2 boxes. The reward structure promotes efficiency and accuracy: +1 for each box correctly placed on a target, -1 for each box off-target, +10 upon full task completion, and -0.1 per action penalty.
Figure 3: Sokoban Environment
Mathematical Problem Solving: Focuses on verifiable mathematical problems. We use the Eurus-2-RL-Math dataset cui2025processreinforcementimplicitrewards for training. Model performance is evaluated on the following test sets: MATH500 hendrycks2021measuringmathematicalproblemsolving , AIME 2024 li2024numinamath , AMC li2024numinamath , AIME 2025, OlympiadBench he2024olympiadbench , and Minerva lewkowycz2022solving .
Knights and Knaves (K&K) Puzzles xie2025memorizationlargelanguagemodels : This dataset consists of algorithmically generated reasoning puzzles. Characters are either Knights (always truthful) or Knaves (always lying). The goal is to determine each character’s type based on their statements.
Hardware.
Verl Framework: Utilizes one machine with 8 $\times$ A100 80GB GPUs for all tasks.
Echo Framework: On the inference side, we use a distributed network of 6 nodes (Parallax ): 3 $\times$ RTX 5090 and 3 $\times$ Mac M4 Pro . On the training side, we use a server equipped with 4 $\times$ A100 GPUs.
Training Recipe.
RL training employs the GRPO algorithm under both the Verl and Echo reinforcement learning frameworks. Hyperparameter settings are task-specific:
Sokoban (Base model: Qwen3-4B): Learning Rate: 1e-6, Batch Size: 6, Rollout N: 16, KL Coefficient: 0.001.
Mathematical Problem Solving (Base model: Qwen2.5-7B): Learning Rate: 5e-7, Batch Size: 256, Rollout N: 8, KL Coefficient: 0.0.
K&K (Base model: Qwen3-32B, trained with LoRA): Learning Rate: 1e-6, Batch Size: 64, Rollout N: 8, KL Coefficient: 0.001, LoRA rank: 64, LoRA alpha: 32.
In the Echo framework, we employ the Sequential mechanism (accuracy-centric) for now, such that we can precisely verify the correctness of the implementation of Echo. Specifically, the model is synchronized in each inference step during rollout generation. This setup ensures the training process logically remains fully aligned with Verl throughout the experiments. The tuning of more-advanced synchronization algorithms is left for future work.
3.2 Results
End-to-End Comparison of Training Convergence.
Figure 2 shows the training curves of three RL tasks specified in the above section. In every plot, we contrast the canonical, fully co-located Verl baseline with our decoupled implementation running on Echo. In the Figure 2(a), success rates fluctuate heavily by design, yet the blue and orange traces track each other throughout training. After around 160 steps, both variants stabilise around a 0.40 success rate, again indicating no degradation in final performance. In the Figure 2(b), the two curves are almost indistinguishable: they accelerate rapidly to 0.8 in the first 80 steps, then converge smoothly toward the 0.9 ceiling. Wall-clock measurements show the same number of actor steps to reach 0.85, validating that our synchronisation protocol maintains sample efficiency. In the Figure 2(c), we scale our experiment model to 32B. Both systems begin around a mean critic reward of 2.3 and rise steadily. From step 10 onwards, the decoupled variant even edges ahead, finishing at $\sim$2.8 versus the baseline’s $\sim$2.75. The overlap in the early phase and the matching plateau confirm that separating rollout generation does not introduce optimization bias.
Performance Improvement by Echo.
In the box-pushing task, the Qwen3-4B model post-trained with Echo(GRPO) surpassed the Qwen3-32B, improving by $4\%$. As shown in Table 1. In math reasoning tasks, we post-trained Qwen2.5-7B with Echo on Eurus-2-RL-Math datasetcui2025processreinforcementimplicitrewards , and it outperformed the Qwen2.5-32B on all six test datasets, achieving a $12\%$ improvement on average, as shown in Table 2. Our experiments on the K&K task show that training under Echo significantly improves model performance, especially in more difficult multi-person scenarios (6-8 people). The Qwen3-32B model fine-tuned with GRPO and LoRA achieved near-perfect scores (all metrics $\geq$ 0.99), surpassing o3-mini-high, DeepSeek-R1 and o4-mini, achieving the best performance. See Table 3 for detailed results.
Table 1: Model performance on Sokoban task
Model Success Rate (%)
Qwen3-4B 21.8
Qwen3-32B 30.0
Qwen3-4B-Echo(GRPO) 34.0
Table 2: Model performance on math reasoning tasks. For AIME and AMC, the results are avg.@32
Model AIME24 AIME25 AMC MATH-500 OlympiadBench Minerva Avg.
Qwen2.5-7B 2.7% 1.9% 22.0% 44.6% 19.7% 20.9% 18.6%
Qwen2.5-32B 5.3% 2.1% 27.9% 62.4% 25.4% 33.5% 26.1%
Qwen2.5-7B-Echo(GRPO) 13.1% 6.9% 45.6% 75.4% 37.0% 50.7% 38.1%
Table 3: Model performance on K&K logic puzzle task across different degrees of difficulty
Model 2 3 4 5 6 7 8
Qwen3-32B 0.98 0.99 0.98 0.99 0.98 0.96 0.95
Deepseek-R1 1.00 0.97 0.95 0.93 0.91 0.93 0.91
o3-mini-high 1.00 1.00 1.00 1.00 0.99 0.98 0.98
o4-mini 1.00 1.00 0.96 0.94 0.97 0.93 0.87
Qwen3-32B-Echo(GRPO w/ Lora) 0.99 1.00 1.00 1.00 0.99 1.00 0.99
4 Conclusion and Future Work
We presented Echo, a dual-swarm RL system that disentangles trajectory sampling from policy optimisation so each phase can exploit the hardware that suits it best. Echo combines (i) a fully decentralised, pipeline-parallel inference engine (Parallax) capable of harnessing heterogeneous edge devices and (ii) an enhanced Verl trainer that retains broad algorithm support and production-grade LoRA fine-tuning. Two lightweight synchronisation protocols—sequential for accuracy and asynchronous for efficiency—bound policy lag without requiring datacentre-class interconnects and can be swapped with only minimal orchestration changes. Experiments on Qwen3-4B, Qwen2.5-7B, and Qwen3-32B show that Echo matches a fully co-located baseline in convergence speed and final reward while off-loading trajectory generation to commodity edge hardware, demonstrating the viability of large-scale RL on geographically dispersed resources.
Future Work.
The dominant cost in our current deployment is model-parameter synchronisation. We plan to attack this bottleneck on two fronts:
  1. 1.
    Reducing synchronisation frequency. We will design runtime-adaptive sync policies that use training-time statistics, e.g., KL divergence between successive policies, gradient-norm trends, or replay-buffer skew, to decide whether and when the inference swarm should refresh its weights, thereby avoiding unnecessary transfers.
  2. 2.
    Reducing synchronisation volume. We intend to explore compression and quantisation techniques tailored to single-direction snapshot distribution, such as low-precision LoRA deltas, sparsity-aware encoding, or entropy-constrained chunking, to shrink each weight push by an order of magnitude.
Together, adaptive frequency control and communication-efficient weight encoding will pave the way for deploying Echo on even wider classes of edge devices, thereby unlocking a truly global, decentralised pool of RL compute.